# Understanding Code Understandability Improvements in Code Reviews

### v1

@@ 3.2 Data Collection
To answer our research questions, we select open-source projects with a code review process active in GitHub, i.e., projects where new pull requests are created frequently. Our focus is on the code review comments made by project contributors regarding the code changes submitted through these pull requests. Furthermore, those projects must be non-trivial software systems, e.g., applications, infrastructure components, libraries, frameworks, etc., with many contributors. We start out by using SEART [48] to search for project candidates because it offers appropriate criteria to search for projects. In SEART, we consider the following criteria: (i) Java projects, (ii) at least ten stars (fixed by SEART), (iii) at least one pull request, (iv) at least ten collaborators, (v) not a fork repository.
@@
We select projects with ten or more contributors because we need a number of contributors that justifies the code review process. These criteria contribute to filter out personal projects [37]. SEART returns a list of 9,735 projects.
@@
Next, we filter the projects to select only those where developers frequently submit their code changes for review by another project contributor. Thus, we select projects with at least one pull request per month in 2020, that includes one or more inline code review comments. These requirements ensure that code reviews are commonplace and happen at the level of source code lines. Given that our search was conducted in October 2021, we focus on pull requests submitted during the year 2020. With this filter, we keep 389 projects.
@@
We further refine this set by removing projects where we can not identify the pull requests that were merged, since merging a pull request is a clear criterion to determine whether said pull request was accepted. Merging can be done manually through git commands on the local repository and then pushing to the remote. Alternatively, it can be done by using the merge button directly on the GitHub platform. Pull requests with automated merging have a clear indication of this event on the pull request’s timeline, with its status updated to ”merged.” However, identifying projects with manual merging can be challenging as it requires the use of heuristics to determine the commit that merged the pull request, as explained by Kalliamvakou et al. [37]. We identify eight projects (e.g., apache/gobblin and openjdk/skara) where the pull requests are not merged automatically and decided to remove them. After performing these filtering steps, we have 381 projects with an active code review process that integrates their code changes in their main branch through Github.
@@
Then, we perform a fine-grained filtering of these 381 projects. We manually analyze each one and evaluate whether it is mainly composed of code. As a result, we exclude eight projects that do not fit our software criteria, such as those classified as documentation. Also, we constrain our investigation to projects where most code review comments are written in English. We remove seven projects where most code review comments are not in English. In addition, one project (returntocorp/semgrep) didn’t primarily use Java, and another (apollographql/apollo-android) was transitioning from Java to Kotlin. As a result, we remove both. Finally, in one project (fisco-bcos/web3sdk) the code review comments are all made by bots, such as “sonarcloud”. This project is also removed, leaving us with a final total of 363 projects. The list can be found at https://codeupcrc.github.io/projects.html. The selected 363 projects comprise more than 349,000 code review comments in total made in 2020.
@@
We download these code review comments using the GitHub GraphQL API  [^7] https://docs.github.com/en/graphql. Since we need to manually evaluate and classify the code review comments, we extract a statistically representative sample. We chose a confidence level of 95% and a margin of error of 2%, resulting in a sample size of 2,401 code review comments for an unlimited population size. We employ a stratified random sampling [49] strategy, as the resulting sample can be more representative of the investigated population of code review comments than if we just employed simple random sampling [50, Chapter 11], where projects with few code review comments might not be represented. We randomly select code review comments from the projects in proportion to the total number of code review comments each project has. This means that we select more code review comments from projects that have a higher number of code review comments. As a result, the minimum number of code review comments per project is 1, the maximum is 15, and the median is 7. The complete list of analyzed comments can be browsed at https://codeupcrc.github.io. Table I presents descriptive statistics of these code review comments and their associated discussions. The table indicates that in some cases, code review comments can be zero characters. These comments comprise only images (Figure 3 presents an example), bullet points without accompanying text, emojis, or are simply empty.
@@ 3.3 Identifying Understandability Improvements
To answer the research questions of this study, we need to classify whether reviewer comments suggest understandability improvements. In this classification process, we need to identify the intention of the reviewer based on their comment. Sometimes, this intention is clear and explicit in the comment text. For example, in Figure 5, the reviewer asked to rename a local variable because “it will help code more readable down into method […]”. However, other comments are terse, and the intent is not explicitly stated. For example, in Figure 6, the reviewer asks to rename deadBrokers to deadBrokersWithUnknownCapacity but the intention is not explicit. Although the reviewer does not write it, the most likely motivation is that the new name helps developers better understand the purpose of this variable.
@@
# chars/discuss 224.2 435.7 0 50 115 247 45,101
# comments/discuss 1.9 1.2 1 1 2 2 43
When the inline comment is not enough, it is necessary to consider more information in the context, i.e., the replies to the comment, the source code, and the patch (when available). We first check whether the suggestion does not change the code’s functionality. For example, in Figure 7, the reviewer asks the author of the pull request to pass another object to function contains instead of the variable eol. We need to evaluate whether the suggestion changes the function of the code snippet to which the suggestion refers because it could indicate that the reviewer is suggesting a correction or different functionality. When we look at the patch, we confirm that the suggestion is behavior-preserving, i.e., the old and the new versions produce the same result given the same input. Thus we consider it as an understandability improvement. However, we have found some comments where the logic of the source code changed, but the function is the same. For instance, in Figure 4, the reviewer suggests changing how a list is processed in the stream, but the expected result of the stream is preserved. We also consider these cases to be potential understandability improvements.
@@
Figure 4: Example of suggested change to preserve the function but change the logic [52].
In summary, we classify a code review comment as suggesting an understandability improvement if it fits within one of the following cases: (i) the intent to improve understandability is explicit in the text; (ii) it is a comment about aspects that do not affect the compilation or the execution of the program, i.e., code comments or code formatting, (iii) the refactoring that it suggests is motivated mainly by improving understandability, as discussed by Fowler [53], and (iv) it does not explicitly state its intent, but it makes a suggestion that is very similar to a comment where the intent is explicit, i.e., the suggestions for improvement require the same change, e.g., in an explicit code review comment, the type Boolean is replaced by boolean to ensure consistency [key=1201], similarly, in an implicit code review comment, the type LinkedList is replaced by ArrayList [key=1089] for the same reason. In a code review comment, the reviewer could suggest an improvement that impacts multiple factors, such as performance and understandability. We also consider those types of improvements that impact understandability. Each comment, including those where the reviewer explicitly mentioned that it pertains to understandability or those that are naturally about understandability (as described in Section 2.2), was classified by at least two authors. Conflicts between answers were resolved through meetings involving at least five authors. In the meetings, we analyzed the conflicts through discussion sessions and resolutions were achieved by consensus. When consensus could not be achieved, the code review comment was discarded and a new one was selected for labeling. In the remainder of this section we explain the types of comments that required additional analysis.
@@
There are scenarios where it is impossible to discern the intention of the reviewer. For example, in a comment [54], the reviewer suggested “better to use putString”. This suggestion can be a related to improving code understandability or improving performance. There is no evidence of any intention, neither in replies nor patches. If we cannot establish intention with some confidence, we discard the comment.
@@
Similarly, in other code review comments, we find tiny comments, e.g., “ditto” [55] and “same here” [56], that do not bring any information about the reviewer’s intention. It would be necessary to look at another code review comment to get the context of the comment, and looking for the original comment could be time-consuming. Since these comments account for approximately 1% of our sample, we have decided to disregard them. However, we have retained them in the dataset labeled as “Discard” for future research.
@@ 4.1 How often do reviewers ask for code understandability improvements in code review comments? (RQ1)
Table II presents the results of this two level classification. Each row indicates the number of comments per scope. The scope of a code review comment can be related to (i) executable source code (Code), i.e., excluding comments and blank lines and focusing on the instructions that will actually be compiled and executed, (ii) high level API documentation (Javadoc), (iii) regular source code comments (Comment), (iv) a string value or expression with strings (String literal), or (v) Others (e.g., annotations). We separated the comments to generate documentation and regular comments because they have different characteristics. Javadoc is well structured, has special annotations to reference specific elements in the documentation, and is used to explain classes, attributes, and methods, often for readers who are not looking at the code. Comment has no prescribed structure and is used to explain the code or the rationale behind it in a more direct manner for readers who are directly looking at a specific part of the code. Also, we analyzed String literals separately because these literals may give developers hints about program behavior, e.g., in log messages, but they are not comments. Finally, annotations, log levels, and other special cases are analyzed separately. We did not consider annotations to be part of the executable source code because, according to the Java Language Specification [^9] https://docs.oracle.com/javase/specs/jls/se12/html/index.html, they have “no effect at run time”.
@@
In the analyzed sample, we found out that 1,012 (42.15%) of the code review comments include suggestions, requests, or discussions about code understandability improvements. This amounts to more than two out of every five code reviews. This result emphasizes that reviewers are worried not only about the correctness and performance of the code but whether the developers are following the best practices of programming and code understandability. We check whether comments suggesting code understandability improvements differ from other comments by comparing their sizes. If one kind of comment tends to be significantly shorter, this may suggest that developers are willing to dedicate less time and effort to it. The 1,012 comments suggesting code understandability improvements tend to be shorter (99.21 characters) than those not related to understandability (117.70 characters). The difference is statistically significant (Mann-Whitney U, $p=0.000001$) but the effect size is negligible (Cliff’s Delta, -0.12). As for the associated source code lines, i.e., the lines where the inline code reviews were added, the mean character counts are 50.52 and 54.15 and the difference is not statistically significant (Mann-Whitney U, $p=0.07$). Overall, we can say that the difference between comments and source lines of code associated to understandability improvements and those that are not is inconsequential.
@@
Considering all code review comments analyzed (2,401), the majority of them (82.2% (1,974 out of 2,401)) are in the scope of Code. Among the 1,974 code review comments whose scope is Code, 33.49% (661 out of 1,974) are related to code understandability improvements. The reviewers in these code review comments have asked code submitters to rename identifiers, remove unnecessary code elements, use alternative code constructs, and others. We discuss the content of those code review comments in Section 4.2. The results support other studies [11, 12] that indicate that improving code quality is one of the main goals of code reviews.
@@
TABLE II: Classification of the 2,401 code review comments in terms of whether they pertain to code understandability (Yes) or not (No). The two groups are further divided based on the scopes to which the code review comments refer (code, JavaDoc, string literals, etc.). The percentages for each row add up to 100%.
@@
Answer to RQ1. We found that 42.15% (1,012/2,401) of the selected discussions are related to code understandability improvements. Among them, 65.31% (661/1,012) are about executable source code, almost 24,01% (157+86 out of 1,012) are related to comments, and more than 10.67% (65+43 out of 1,012) are related to literal strings, annotations, and others. Approximately one out of every three code review comments that refers to executable source code is about understandability. Furthermore, four out of five code review comments that refer to elements other than executable code in a program source file pertain to understandability improvements. Implications. In previous research  [27, 28], the researchers identified ways to improve code understandability at the level of files by analyzing commit messages. Our findings suggest that researchers and developers can also learn from code reviews to improve code understandability at the level of source code snippets. Moreover, our results demonstrate that, automated tools that enhance code understandability can have significant impact in reducing developers’ effort during code reviews. Additionally, our work makes it clear that improving code understandability requires tools that work with both programming language and natural language elements.
@@ 4.2 What are the main issues pertaining to code understandability in code review? (RQ2)
Another understandability smell is complex, long, or inadequate logic. The reviewers pointed out concerns about complex and long alternatives to writing expressions, code constructs, and other statements. For instance, the boolean expression map.keySet().size() > 0 could be replaced by an unique method invocation, !map.isEmpty(). We find this category in 18.2% (70/385) of the observations in our sample. The reviewers have identified redundant method calls (e.g., “Looks like it could be simplified” [key=83]); and long method implementation (e.g., “Maybe you can refactor this logic into a different method, it makes this method quite large…” [key=1413]). We also found comments suggesting that anonymous inner classes are too verbose (e.g., “With a lambda, this would be slightly shorter” [key=58]); and complex if-else implementation with repeated code (“I think a better way to write this method is: [duplicate code within if-else out of if-else statement]” [key=1878]). Moreover, reviewers indicated concerns related to using API classes, such as using a class as type instead of its primitive type (e.g., “Why Boolean and not boolean?” [key=1201]). This smell is more frequent in Call (27), Methods (eleven in the body and two in the declaration clause) and Conditionals (nine).
TABLE III: The frequency of understandability smells (what) found by reviewers and where they are in the source code. The understandability smells were found in declaration, access, or implementation of kind of code elements with an asterisks.
@@
Answer to RQ2. In our representative sample of 385 code reviews, we found eight categories of understandability smells. Missing or inadequate code documentation (86), Bad identifier (78), and Complex, long, or inadequate logic (70) comprise more than 60% of the understandability smells in our sample. They are related to 16 code element types in source code. The three most frequent places, i.e., Method (130), Call (55), and Literal (40), account for more than 58% of the code reviews. Implications. We have discovered a variety of issues affecting code understandability and grouped them into understandability smells. Automated solutions to improve code understandability must deal with these different kinds of issues, with varying levels of context dependence, and tackling multiple modalities: source code as well as unstructured and structured natural language. The results suggest that tools based on LLMs have to be trained on these different languages, as some smells manifest only in Java source code (e.g., Unnecessary code) whereas others occur in natural language text or identifiers, e.g., Bad identifier and Incomplete or inadequate code documentation.

### v2

@@ 3.2 Data Collection
To answer our research questions, we selected open-source projects with a code review process active in GitHub, i.e., projects where new pull requests are created frequently. Our focus is on the code review comments made by project contributors regarding the code changes submitted through these pull requests. Furthermore, those projects must be non-trivial software systems, e.g., applications, infrastructure components, libraries, frameworks, with many contributors.
@@
We started out by using SEART [48] to search for project candidates because it offers appropriate search criteria. In SEART, we considered the following criteria: (i) Java projects, (ii) at least ten stars (fixed by SEART), (iii) at least one pull request, (iv) at least ten contributors, and (v) not a fork repository. We selected projects with ten or more contributors because we need a number of contributors that justifies the code review process. These criteria contribute to filtering out personal projects [37]. SEART returned a list of 9,735 projects.
@@
Next, we filtered the projects to select only those where code changes submitted by developers are regularly reviewed by other contributors. Thus, we selected projects with at least one pull request per month in 2020 that include one or more inline code review comments. These requirements ensure that code reviews are commonplace and happen at the level of source code lines. Given that our search was conducted in October 2021, we focused on pull requests submitted during the year 2020. With this filter, we kept 389 projects.
@@
We further refined this set by removing projects where we can not identify the pull requests that were merged since merging a pull request is a clear criterion to determine whether said pull request was accepted. Merging can be done manually through git commands on the local repository and then pushing to the remote. Alternatively, it can be done by using the merge button directly on the GitHub platform. Pull requests with automated merging have a clear indication of this event on the pull request’s timeline, with its status updated to ”merged.” However, identifying projects with manual merging can be challenging as it requires the use of heuristics to determine the commit that merged the pull request, as explained by Kalliamvakou et al. [37]. We identified eight projects (e.g., apache/gobblin and openjdk/skara) where the pull requests are not merged automatically and decided to remove them. After performing these filtering steps, we kept 381 projects with an active code review process that integrates their code changes in their main branch through GitHub.
@@
Then, we performed a fine-grained filtering of these 381 projects. We manually analyzed each one and evaluated whether it is mainly composed of code. As a result, we excluded eight projects that do not fit our software criteria, such as those classified as documentation. Also, we constrained our investigation to projects where most code review comments are written in English. We removed seven projects where most code review comments are not in English. In addition, one project (returntocorp/semgrep) did not primarily use Java, and another (apollographql/apollo-android) was transitioning from Java to Kotlin. As a result, we removed both. Finally, the code review comments in one project (fisco-bcos/web3sdk) are all made by bots, such as “sonarcloud”. This project was also removed, leaving us with a final total of 363 projects [^6] The list of selected projects can be found at https://codeupcrc.github.io/projects.html.
@@
The selected 363 projects comprise more than 349,000 code review comments made in 2020. We downloaded these code review comments using the GitHub GraphQL API [^7] https://docs.github.com/en/graphql. Since we need to manually evaluate and classify the code review comments, we extracted a statistically representative sample. We chose a confidence level of 95% and a margin of error of 2%, resulting in a sample size of 2,401 code review comments for an unlimited population size. We employed a stratified random sampling [49] strategy, as the resulting sample can be more representative of the investigated population of code review comments than if we just employed simple random sampling [50, Chapter 11], where projects with few code review comments might not be represented. We randomly selected code review comments from the projects in proportion to the total number of code review comments each project has. As a result, the minimum number of code review comments per project is 1, the maximum is 15, and the median is 7. The complete list of analyzed comments can be browsed at https://codeupcrc.github.io. Table I presents descriptive statistics of these code review comments and their associated discussions. The table indicates that, in some cases, code review comments can be zero characters. These comments comprise only images (Figure 3 presents an example), bullet points without accompanying text, emojis, or are simply empty.
@@ 3.3 Identifying Understandability Improvements
To answer the research questions of this study, we first classified whether code review comments suggest understandability improvements or not. In this classification process, we need to identify the intention of the reviewer based on their comment. Sometimes, this intention is clear and explicit in the comment text. For example, in Figure 5, the reviewer asked to rename a local variable because “it will help code more readable down into method […]”. However, other comments are terse, and the intent is not explicitly stated. For example, in Figure 6, the reviewer asks to rename deadBrokers to deadBrokersWithUnknownCapacity but the intention is not explicit. Although the reviewer does not write it, the most likely motivation is that the new name helps developers better understand the purpose of this variable.
@@
# chars/discussion 224.2 435.7 0 50 115 247 45,101
# comments/discussion 1.9 1.2 1 1 2 2 43
When the inline comment is not enough, it is necessary to consider more information in the context, i.e., the replies to the comment, the source code, and the patch (when available). We first checked whether the suggestion does not change the functionality of the code. For example, in Figure 7, the reviewer asks the author of the pull request to pass another object to the function contains instead of the variable eol. We first evaluated whether the suggestion changes the function of the code snippet to which the suggestion refers because it could indicate that the reviewer is suggesting a correction or different functionality. When we look at the patch, we confirm that the suggestion is behavior-preserving, i.e., the old and the new versions produce the same result given the same input. Thus we consider it as an understandability improvement. However, we found some comments where the logic of the source code changed, but the function is the same. For instance, in Figure 4, the reviewer suggests changing how a list is processed in the stream, but the expected result of the stream is preserved. We also consider these cases to be potential understandability improvements.
@@
Figure 4: Example of a suggested change to preserve the function but change the logic [52].
In summary, we classified a code review comment as suggesting an understandability improvement if it fits within one of the following cases: (i) the intent to improve understandability is explicit in the text, (ii) it is a comment about aspects that do not affect the compilation or the execution of the program, i.e., code comments or code formatting, (iii) the refactoring that it suggests is motivated mainly by improving understandability, as discussed by Fowler [53], or (iv) it does not explicitly state its intent, but it makes a suggestion that is very similar to a comment where the intent is explicit, i.e., the suggestions for improvement require the same change, e.g., in an explicit code review comment, the type Boolean is replaced by boolean to ensure consistency [key=1201], similarly, in an implicit code review comment, the type LinkedList is replaced by ArrayList [key=1089] for the same reason. In a code review comment, the reviewer could suggest an improvement that impacts multiple factors, such as performance and understandability. We also consider those types of improvements that impact understandability. Each comment, including those where the reviewer explicitly mentioned that it pertains to understandability or those that are naturally about understandability (as described in Section 2.2), was classified by at least two authors. Conflicts between answers were resolved through meetings involving at least five authors. In the meetings, we analyzed the conflicts through discussion sessions and resolutions were achieved by consensus. When consensus could not be achieved, the code review comment was discarded and a new one was selected for labeling. In the remainder of this section, we explain the types of comments that required additional analysis.
@@
There are scenarios where it is impossible to discern the intention of the reviewer. For example, in a comment [54], the reviewer suggested “better to use putString”. This suggestion can be related to improving code understandability or improving performance. There is no evidence of any intention, neither in replies nor patches. If we cannot establish intention with some confidence, we discard the comment.
@@
Similarly, in other code review comments, we find tiny comments, e.g., “ditto” [55] and “same here” [56], that do not bring any information about the reviewer’s intention. It would be necessary to look at another code review comment to get the context of the comment, and looking for the original comment could be time-consuming. Since these comments account for approximately 1% of our sample, we decided to disregard them. However, we kept them in the dataset labeled as “Discard” for future research.
@@ 4.1 How often do reviewers ask for code understandability improvements in code review comments? (RQ1)
Table II presents the results of our manual classification of code review comments in terms of whether they pertain to code understandability or not. Each row indicates the number of comments per scope, which can be related to (i) executable source code (Code), i.e., excluding comments and blank lines and focusing on the instructions that will actually be compiled and executed, (ii) high-level API documentation (Javadoc), (iii) regular source code comments (Comment), (iv) a string value or expression with strings (String literal), or (v) Others (e.g., annotations). We separated the comments to generate documentation and regular comments because they have different characteristics. Javadoc is well structured, has special annotations to reference specific elements in the documentation, and is used to explain classes, attributes, and methods, often for readers who are not looking at the code. Comment has no prescribed structure and is used to explain the code or the rationale behind it in a more direct manner for readers who are directly looking at a specific part of the code. Also, we analyzed String literals separately because these literals may give developers hints about program behavior, e.g., in log messages, but they are not comments. Finally, annotations, log levels, and other special cases are analyzed separately. We did not consider annotations to be part of the executable source code because, according to the Java Language Specification [^9] https://docs.oracle.com/javase/specs/jls/se12/html/index.html, they have “no effect at run time”.
@@
In the analyzed sample, we found out that 1,012 (42.15%) of the code review comments include suggestions, requests, or discussions about code understandability improvements. This amounts to more than two out of every five code reviews. This result emphasizes that reviewers are worried not only about the correctness and performance of the code but also whether the developers are following the best practices of programming and code understandability. We check whether comments suggesting code understandability improvements differ from other comments by comparing their sizes. If one kind of comment tends to be significantly shorter, this may suggest that developers are willing to dedicate less time and effort to it. The 1,012 comments suggesting code understandability improvements tend to be shorter (99.21 characters) than those not related to understandability (117.70 characters). The difference is statistically significant (Mann-Whitney U, $p=0.000001$) but the effect size is negligible (Cliff’s Delta, -0.12). As for the associated source code lines, i.e., the lines where the inline code reviews were added, the mean character counts are 50.52 and 54.15 and the difference is not statistically significant (Mann-Whitney U, $p=0.07$). Overall, we can say that the difference between comments and source lines of code associated with understandability improvements and those that are not is inconsequential.
@@
Considering all code review comments analyzed (2,401), the majority of them (82.2%, 1,974 out of 2,401) are in the scope of Code. Among the 1,974 code review comments whose scope is Code, 33.49% (661 out of 1,974) are related to code understandability improvements. The reviewers in these code review comments have asked code submitters to rename identifiers, remove unnecessary code elements, use alternative code constructs, and others. We discuss the content of those code review comments in Section 4.2. The results support other studies [11, 12] that indicate that improving code quality is one of the main goals of code reviews.
@@
TABLE II: Classification of the 2,401 code review comments in terms of whether they pertain to code understandability or not. The two groups are further divided based on the scopes to which the code review comments refer. The percentages for each row add up to 100%.
@@
Answer to RQ1. We found that 42.15% (1,012/2,401) of the selected discussions are related to code understandability improvements. Among them, 65.31% (661/1,012) are about executable source code, almost 24,01% (157+86 out of 1,012) are related to comments, and more than 10.67% (65+43 out of 1,012) are related to literal strings, annotations, and others. Approximately one out of every three code review comments that refer to executable source code is about understandability. Furthermore, four out of five code review comments that refer to elements other than executable code in a program source file pertain to understandability improvements. Implications. In previous research  [27, 28], the researchers identified ways to improve code understandability at the level of files by analyzing commit messages. Our findings suggest that researchers and developers can also learn from code reviews to improve code understandability at the level of source code snippets. Moreover, our results demonstrate that automated tools that enhance code understandability can have a significant impact in reducing developers’ effort during code reviews. Additionally, our work makes it clear that improving code understandability requires tools that work with both programming language and natural language elements.
@@ 4.2 What are the main issues pertaining to code understandability in code review? (RQ2)
Another understandability smell is complex, long, or inadequate logic. The reviewers pointed out concerns about complex and long alternatives to writing expressions, code constructs, and other statements. For instance, the boolean expression map.keySet().size() > 0 could be replaced by a unique method invocation, !map.isEmpty(). We find this category in 18.2% (70/385) of the observations in our sample. The reviewers have identified redundant method calls (e.g., “Looks like it could be simplified” [key=83]); and long method implementation (e.g., “Maybe you can refactor this logic into a different method, it makes this method quite large…” [key=1413]). We also found comments suggesting that anonymous inner classes are too verbose (e.g., “With a lambda, this would be slightly shorter” [key=58]); and complex if-else implementation with repeated code (“I think a better way to write this method is: [duplicate code within if-else out of if-else statement]” [key=1878]). Moreover, reviewers indicated concerns related to using API classes, such as using a class as type instead of its primitive type (e.g., “Why Boolean and not boolean?” [key=1201]). This smell is more frequent in Call (27), Methods (eleven in the body and two in the declaration clause), and Conditionals (nine).
TABLE III: The frequency of understandability smells found by reviewers and where they are in the source code. For the code elements marked with an asterisk, the understandability smells were found in their declaration, access, or implementation.
@@
Answer to RQ2. In our representative sample of 385 code reviews, we found eight categories of understandability smells. Missing or inadequate code documentation (86), Bad identifier (78), and Complex, long, or inadequate logic (70) comprise more than 60% of the understandability smells in our sample. They are related to 16 source code element types. The three most frequent places, i.e., Method (130), Call (55), and Literal (40), account for more than 58% of the code reviews. Implications. We discovered a variety of issues affecting code understandability and grouped them into understandability smells. Automated solutions to improve code understandability must deal with these different kinds of issues, with varying levels of context dependence, and tackling multiple modalities: source code as well as unstructured and structured natural language. The results suggest that tools based on LLMs have to be trained in these different languages, as some smells manifest only in Java source code (e.g., Unnecessary code) whereas others occur in natural language text or identifiers, e.g., Bad identifier and Incomplete or inadequate code documentation.